# QKD security proofs for decoy-state BB84: protocol variations, proof techniques, gaps and limitations

### v1

@@ 4.1 Source-replacement scheme
Broadly, the source-replacement scheme [22, 23] is used, as the name suggests, to replace the actual source with a virtual entangled source, which when coupled with a measurement on Alice’s subsystem can mimic the actual source. The exact entangled source and measurement used depends on the proof-technique — some proof techniques (phase error and EUR) use this to reduce Alice’s subsystem to a qubit system to simplify the analysis, while others (postselection and EAT-based) instead require an IID source preparation that need not be a qubit.
@@ 4.1.1 Shield system
4.2 2-universal hashing
Informally, hash functions usually serve the purpose of “compressing” long input strings into shorter output strings, in such a way that the event of different input values giving the same output value (usually called a collision) is “rare” in some sense. 2-universal hashing serves as a rigorous way to formalize this intuition, as seen from its definition: [^16] There exists a generalization of 2-universal hashing referred to as $\delta$-almost-2-universal hashing [25], but we will not cover it in detail within this work.
@@
Given some domain $D$ and finite codomain $D^{\prime}$, a 2-universal hash family is a set $\mathcal{H}$ consisting of functions $h:D\to D^{\prime}$, together with a probability distribution over $\mathcal{H}$, such that if we randomly draw a function $h$ from $\mathcal{H}$ according to that distribution, we have:
$\text{For all }x,y\in D\text{ such that }x\neq y,\quad\Pr[h(x)=h(y)]\leq\frac{1}{|D^{\prime}|},$ (8)
where the probability is evaluated with respect to the distribution over $\mathcal{H}$.
@@
We strongly emphasize that 2-universal hashing is a very distinct concept from using fixed choices of computationally-hard functions (or functions with “heuristic hardness estimates”, such as SHA-3), and should never be confused with those notions. In particular, we stress that no computational-hardness assumptions or heuristics are needed to construct 2-universal hash families that satisfy the above definition. [^17] To achieve this, it was critical that the function $h$ is drawn randomly from some family in 2-universal hashing — notice that if $h$ had been a fixed function and $|D|>|D^{\prime}|$, then it would have been impossible to find a bound of the form in Eq. 8 without further assumptions, since we would necessarily have a collision by the pigeonhole principle. Hence 2-universal hashing allows us to rigorously ensure information-theoretic security, without computational assumptions.
Slightly informally, we often refer to the overall process of randomly drawing a function $h$ from a 2-universal hash family and applying $h$ to some classical register as “applying a 2-universal hash”, and we will refer to the output value as the hash. For most 2-universal hash families used in QKD, the relevant distribution over $\mathcal{H}$ is simply a uniform distribution, though this is not a required property in the definition. Furthermore, the codomain $D^{\prime}$ usually just consists of bitstrings of some length $\ell_{\mathrm{hash}}$, so the above bound reduces to
$\text{For all }x,y\in D\text{ such that }x\neq y,\quad\Pr[h(x)=h(y)]\leq\frac{1}{2^{\ell_{\mathrm{hash}}}},$ (9)
i.e. it is exponentially small with respect to the length $\ell_{\mathrm{hash}}$ of the output strings. We now highlight some other important properties of 2-universal hashing.
@@
There exist explicit constructions of 2-universal hash families that are reasonably efficient, with respect to both computational complexity and the randomness required to draw from that family. (We again stress that no computational-hardness assumptions are needed to rigorously prove that these constructions satisfy the defining property, Eq. 8.) For instance, when the domain and codomain are both bitstrings of fixed-lengths, a commonly used choice for a 2-universal hash family is the set of Toeplitz matrices, as described in Ref. [26] (see also Ref. [27] for a discussion specific to Toeplitz hashing). This approach offers the advantage that there exist algorithms for efficiently computing the output of a Toeplitz matrix applied to a bitstring, and the number of random bits required to choose such a matrix is linear with respect to the lengths of the input and output strings.
@@
As noted in Section 3, 2-universal hashing is used in many aspects of QKD, for instance in the following steps.
  • Most straightforwardly, it is used in the error verification step, where a 2-universal hash is used to attempt to check whether Alice and Bob obtained matching pre-amplification strings. The fact that it ensures a rigorous bound on the “collision probability” is highly relevant in analyzing this step, as we describe in Section 4.3.
  • It is also applied in many commonly used procedures for privacy amplification. Here, the bound on the collision probability is less “directly” relevant; the key property that ensures security is the Leftover Hashing Lemma (Section 4.4).
  • A slightly strengthened version of 2-universal hashing can also serve as a possible method for the construction of the authenticated classical channel, via Wegman-Carter authentication [7, 8], though we do not consider this within the scope of the QKD security proof itself.
@@ 4.3 The correctness condition
By definition, the correctness condition requires us to show that the probability that the final keys differ and the protocol accepts (or that it produces a nonzero key length, in the variable-length case) is small. We first note that for protocols with error verification, this event is a “stricter” event than the event of the pre-amplification strings differing and the error verification step accepting. [^18] This is because the final keys can only differ if the pre-amplification strings differ, and the protocol can only accept (or produce a nonzero final key length) if the error verification step accepts. Thus it suffices to upper-bound the probability of the latter event. This turns out to be fairly straightforward for error verification based on 2-universal hashing: observe that if the error verification hash is a bitstring of length $\ell_{\mathrm{hash}}$, then
[…]
$\leq$ $2^{\ell_{\mathrm{hash}}},$ (10)
@@
where the second line is Bayes’ rule, the third line holds simply because probabilities are upper bounded by $1$, and the fourth line follows from the defining property of 2-universal hashing (Definition 4.1). Hence by choosing $\ell_{\mathrm{hash}}$ appropriately, the correctness condition is satisfied.
@@
An important aspect of this argument is that it relies solely on the properties of error verification based on 2-universal hashing; in particular, it has no dependence at all on any properties of the preceding error correction step. This implies that for protocols using such an error verification procedure, the error correction procedure can be designed such that it is only guaranteed to succeed with high probability on the honest behaviour (so the completeness condition is fulfilled), rather than having to consider any form of dishonest behaviour at that step, since the correctness condition is ensured by the error verification step alone. Hence this form of error verification usually simplifies the security analysis. Note however that this is not the case for some of the variants discussed in Section 3.4.2; a more elaborate analysis may be required for some of those cases.
@@
We highlight that the correctness condition does not require one to show that conditioned on the protocol accepting, the probability that the final keys differ is small — it only requires us to bound the probability of the joint event, rather than this conditional probability. In fact, one can construct explicit examples of attacks by Eve in which conditioned on the protocol accepting, the probability that the final keys differ is still large, i.e.
[…]
@@
Therefore, it would be a misconception to claim that this conditional probability is small for QKD protocols satisfying the security definitions. However, as discussed in Refs. [5, 6], the composable security framework ensures that the correctness condition in the stated form is already enough to certify the desired composable properties of the protocol, despite the lack of a bound on this conditional probability.
@@ 4.4 Leftover Hashing Lemma
A critical result for proving secrecy in a QKD security proof is the Leftover Hashing Lemma [19, 4, 28] — while there are many variants, we focus on the one in [4, Proposition 9] for the purposes of this discussion, as it is most compatible with the definitions we introduced in Section 2. We can summarize the lemma as follows: for any state $\rho_{CQ}$ with classical $C$, if we apply a 2-universal hash to $C$ such that the output is a bitstring $K$ of length $\ell$, and denote the resulting state as $\rho_{KQH}$ where $H$ is a register storing the choice of hash function, then [^19] In Ref. [4] (Proposition 9), a slightly generalized version of this lemma was presented for subnormalized states.
$d\left(\rho_{KQH},\omega^{\ell}_{K}\otimes\rho_{QH}\right)\leq 2^{-\frac{1}{2}(H_{\mathrm{min}}^{\overline{\varepsilon}}(C|Q)_{\rho}-\ell+2)}+2\overline{\varepsilon},$ (12)
@@
As noted in Section 3.5, there exist privacy amplification procedures not based on 2-universal hashing, such as Trevisan’s extractor [29]. However, the qualitative properties of those procedures are mostly similar to those of the Leftover Hashing Lemma. Furthermore, as mentioned there, additional care is necessary when analyzing privacy amplification if the pre-amplification strings are not bitstrings of fixed-length, since this may not technically fit the conditions of the Leftover Hashing Lemma. We briefly discuss this in Section 5.6.2 later, and a full rigorous treatment of this issue can be found in Ref. [30].
@@ 4.6.1 Source maps
Formally, given the set of (infinite-dimensional) states $\{\rho_{i}\}$ used in the protocol, we find a set of (finite-dimensional) states $\{\tau_{i}\}$ such that there exists a channel $\Phi$ where $\rho_{i}=\Phi(\tau_{i})$ for all $i$. This channel $\Phi$ constitutes the source map which is ‘given’ to Eve, and thus any security analysis that assumes that Alice prepares the states $\{\tau_{i}\}$ would also hold for the actual protocol where Alice prepares the states $\{\rho_{i}\}$ as this only gives Eve more power. See [11, Lemma 8] for a more formal proof of this.
@@ 4.6.2 Squashing maps
Simple squashers [33, 37] Protocol-dependent construction — need to check for each protocol separately (e.g. - equal loss in all detectors for active BB84 setup, no dark counts.) Proof-technique independent [11, Lemma 6]
@@
Flag-state squasher [38] Block-diagonal structure of POVM elements EAT-based proof techniques [31, Appendix B], IID analysis [38, Theorem 1]
Dimension-reduction method [39] None EAT-based proof techniques [^27] Note that although this has been used for CV-QKD protocols, the extension to DV-QKD protocols is straightforward. [41, Theorem 2], IID analysis [39]
Weight-preserving flag-state squasher [11, Section IV.B.1.] Block-diagonal structure of POVM elements Proof-technique independent [11, Lemma 6]
Proof-technique independent squashers: The simple squashers and weight-preserving flag-state squasher use the same general principle [34, 33, 35, 37] depicted in Fig. 2, similar to source maps. These squashers allow us to directly replace the infinite-dimensional POVM with a finite-dimensional POVM for a QKD security proof. That is, they work by proving that the security of the QKD protocol with the finite-dimensional POVM implies the security of the QKD protocol with the infinite-dimensional POVM [11, Lemma 6]. The security proof for the former protocol can then be obtained via any proof technique of one’s choice.
@@ 5.1.2 Iterative sifting
Protocols with on-the-fly announcements (but with a fixed number of total rounds) are easily handled in the GEAT framework [46], but then suffer from the sequential assumptions. For techniques to additionally handle the scenario where the total number of rounds also depends on these announcements, see Ref. [47] for methods in phase error correction (Section 6.2) and EUR (Section 6.1) frameworks.
@@ 5.3.1 Collective attacks are not a good approximation to practical restrictions
A common assumption made in some existing work is that Eve’s attack is IID, i.e., she interacts with Alice’s signal states in each round exactly the same way, independent of prior protocol rounds. We emphasise here that this assumption cannot be motivated by limiting Eve to currently practical attacks. That is, with currently available technology, it is easy to implement a non-IID attack.
@@ 5.3.3 Use of concentration inequalities
Hoeffding’s Inequality [49, Theorem 2] $X_{1},\dots,X_{n}$ independent random variables with $a_{i}\leq X_{i}\leq b_{i}$. $\Pr[\absolutevalue{\sum_{i}(X_{i}-E[X_{i}])}\geq t]\leq 2\exp(-\frac{2t^{2}}{\sum_{i}^{n}\left(b_{i}-a_{i}\right)^{2}})$
@@
Serfling inequality [50, Corollary 1.1] $X_{1},\dots,X_{n}$ drawn randomly without replacement from values $x_{1},\dots,x_{N}$ for $n\leq N$ $\Pr[\sum_{i}X_{i}-n\mu\geq nt]\leq 2\exp(-\frac{2nt^{2}}{(1-f_{n}^{*})(b-a)})$, for $\mu=\sum_{i}x_{i}/n,f_{n}^{*}=\frac{n-1}{N},a=\min_{i}x_{i}$ and $b=\max_{i}x_{i}$.

### v2

@@ 4.1 Source-replacement scheme
Broadly, the source-replacement scheme [12, 26] is used, as the name suggests, to replace the actual source with a virtual entangled source, which when coupled with a measurement on Alice’s subsystem can mimic the actual source. The exact entangled source and measurement used depends on the proof-technique — some proof techniques (phase error and EUR) use this to reduce Alice’s subsystem to a qubit system to simplify the analysis, while others (postselection and EAT-based) instead require an IID (independent and identically distributed) source preparation that need not be a qubit.
@@ 4.1.1 Shield system
4.2 Universal and almost-universal hashing
Informally, hash functions usually serve the purpose of “compressing” long input strings into shorter output strings, in such a way that the event of different input values giving the same output value (usually called a collision) is “rare” in some sense. Universal and almost-universal hashing procedures serve as a rigorous way to formalize some aspects of this intuition, as seen from the definition [18, 98]:
@@ 4.2 Universal and almost-universal hashing
Given some domain $D$ and finite codomain $D^{\prime}$, a $\delta$-almost-universal2 ($\delta$-AU2) hashing procedure consists of a set $\mathcal{H}$ of functions $h:D\to D^{\prime}$, together with a specific probability distribution $P_{\mathrm{HASH}}$ over $\mathcal{H}$, such that if we randomly draw a function $h$ according to the distribution $P_{\mathrm{HASH}}$, we have:
For all $\text{For all }x,y\in D\text{ such that }x\neq y,\quad$ (8)
$\Pr[h(x)=h(y)]\leq\delta,$ (9)
where the probability is evaluated with respect to the distribution $P_{\mathrm{HASH}}$. A special case of $\delta$-almost-universal2 hashing is universal2 hashing, which simply refers to the case where
$\delta=\frac{1}{|D^{\prime}|}.$ (10)
@@
We strongly emphasize that $\delta$-AU2 hashing is a very distinct concept from using fixed choices of computationally-hard functions (or functions with “heuristic hardness estimates”, such as SHA-3), and should never be confused with those notions. In particular, we stress that no computational-hardness assumptions or heuristics are needed to construct $\delta$-AU2 hashing procedures that satisfy the above definition. [^19] To achieve this, it was critical that the function $h$ is drawn randomly in $\delta$-AU2 hashing — notice that if $h$ had been a fixed function and $|D|>|D^{\prime}|$, then it would have been impossible to find a bound of the form in Eq. 8 without further assumptions, since we would necessarily have a collision by the pigeonhole principle. Hence $\delta$-AU2 hashing allows us to rigorously ensure information-theoretic security, without computational assumptions.
Slightly informally, we often refer to the overall process of randomly drawing a function $h$ according to $P_{\mathrm{HASH}}$ and applying $h$ to some classical register as “applying a $\delta$-AU2 hash”, and we will refer to the output value as the hash. For most $\delta$-AU2 hashing procedures used in QKD, the distribution $P_{\mathrm{HASH}}$ is simply a uniform distribution, though the definition we considered does not specifically require this distribution to be uniform. Furthermore, the codomain $D^{\prime}$ usually just consists of bitstrings of some length $\ell_{\mathrm{hash}}$, in which case the bound for universal2 hashing reduces to
For all $\text{For all }x,y\in D\text{ such that }x\neq y,\quad$ (11)
$\Pr[h(x)=h(y)]\leq\frac{1}{2^{\ell_{\mathrm{hash}}}},$ (12)
i.e. it is exponentially small with respect to the length $\ell_{\mathrm{hash}}$ of the output strings. We now highlight some other important properties of $\delta$-AU2 hashing.
@@
There exist explicit constructions of $\delta$-AU2 hashing procedures that are reasonably efficient, with respect to both computational complexity and the randomness required to sample according to $P_{\mathrm{HASH}}$. (We again stress that no computational-hardness assumptions are needed to rigorously prove that these constructions satisfy the defining property, Eq. 8.) For instance, when the domain and codomain are both bitstrings of fixed lengths, a commonly used choice for a universal2 hashing procedure is to sample uniformly from the set of Toeplitz matrices, as described in Ref. [65] (see also Ref. [54] for a discussion specific to Toeplitz hashing). This approach offers the advantage that there exist algorithms for efficiently computing the output of a Toeplitz matrix applied to a bitstring, and the number of random bits required to choose such a matrix is linear with respect to the lengths of the input and output strings.
@@
As noted in Section 3, $\delta$-AU2 hashing is used in many aspects of QKD, for instance in the following steps.
  • Most straightforwardly, it is used in the error verification step, where a $\delta$-AU2 hash is used to attempt to check whether Alice and Bob obtained matching pre-amplification strings. The fact that it ensures a rigorous bound on the “collision probability” is highly relevant in analyzing this step, as we describe in Section 4.3.
  • Universal2 hashing is also applied in many commonly used procedures for privacy amplification. Here, the bound on the collision probability is less “directly” relevant; the key property that ensures security is the Leftover Hashing Lemma (Section 4.4), though the proof of that lemma relies on that bound in an intermediate step.
  • A slightly strengthened version of $\delta$-AU2 hashing can also serve as a possible method for the construction of the authenticated classical channel, via Wegman-Carter authentication [118, 81], though we do not consider this within the scope of the QKD security proof itself.
@@ 4.3 The correctness condition
By definition, the correctness condition requires us to show that the probability that the final keys differ and the protocol accepts (or that it produces a nonzero key length, in the variable-length case) is small. We first note that for protocols with error verification, this event is a “stricter” event than the event of the pre-amplification strings differing (which we denote via $\Omega_{\mathrm{diff}}$) and the error verification step accepting (which we denote via $\Omega_{\mathrm{EV}}$). [^20] This is because the final keys can only differ if the pre-amplification strings differ, and the protocol can only accept (or produce a nonzero final key length) if the error verification step accepts. Thus it suffices to upper-bound the probability of the latter event. This turns out to be fairly straightforward for error verification based on $\delta$-AU2 hashing: observe that this gives
$\Pr[\Omega_{\mathrm{diff}}\text{ AND }\Omega_{\mathrm{EV}}]$
$=$ $\Pr[\Omega_{\mathrm{EV}}\;|\;\Omega_{\mathrm{diff}}]\Pr[\Omega_{\mathrm{diff}}]$
$\leq$ $\Pr[\Omega_{\mathrm{EV}}\;|\;\Omega_{\mathrm{diff}}]$
$\leq$ $\delta,$ (13)
@@
where the third line holds simply because probabilities are upper bounded by $1$, and the fourth line follows from the defining property of $\delta$-AU2 hashing (Definition 4.1). Hence the correctness condition is satisfied with $\varepsilon^{\mathrm{correct}}=\delta$.
@@
An important aspect of this argument is that it relies solely on the properties of error verification based on $\delta$-AU2 hashing; in particular, it has no dependence at all on any properties of the preceding error correction step. This implies that for protocols using such an error verification procedure, the error correction procedure can be designed such that it is only guaranteed to succeed with high probability on the honest behaviour (so the completeness condition is fulfilled), rather than having to consider any form of dishonest behaviour at that step, since the correctness condition is ensured by the error verification step alone. Hence this form of error verification usually simplifies the security analysis. Note however that this is not the case for some of the variants discussed in Section 3.4.2; a more elaborate analysis may be required for some of those cases.
@@
We highlight that the correctness condition does not require one to show that conditioned on the protocol accepting, the probability that the final keys differ is small — it only requires us to bound the probability of the joint event, rather than this conditional probability. In fact, one can construct explicit examples of attacks by Eve in which conditioned on the protocol accepting, the probability that the pre-amplification strings differ is still large, i.e.
$\Pr[\Omega_{\mathrm{diff}}\;|\;\Omega_{\mathrm{EV}}]\approx 1,$ (14)
@@
as we show in Section 5.4.6. Therefore, it would be a misconception to claim that the correctness condition ensures this probability is small. However, as discussed in Refs. [89, 82], the composable security framework ensures that the correctness condition in the stated form is already enough to certify the desired composable properties of the protocol, despite the lack of a bound on this conditional probability.
@@ 4.4 Leftover Hashing Lemma
A critical result for proving secrecy in a QKD security proof is the Leftover Hashing Lemma [85, 105, 29] — while there are many variants, we focus on the one in [105, Proposition 9] for the purposes of this discussion, as it is most compatible with the definitions we introduced in Section 2. We can summarize the lemma as follows: for any state $\rho_{CQ}$ with classical $C$, if we apply a universal2 hash to $C$ such that the output is a bitstring $K$ of length $\ell$, and denote the resulting state as $\rho_{KQH}$ where $H$ is a register storing the choice of hash function, then [^21] In Ref. [105] (Proposition 9), a slightly generalized version of this lemma was presented for subnormalized states. There also exist versions of this lemma based on $\delta$-AU2 hashing rather than universal2 hashing.
$d\left(\rho_{KQH},\omega^{\ell}_{K}\otimes\rho_{QH}\right)\leq\frac{1}{2}2^{-\frac{1}{2}(H_{\mathrm{min}}^{\overline{\varepsilon}}(C|Q)_{\rho}-\ell+2)}+2\overline{\varepsilon},$ (15)
@@
As noted in Section 3.5, there exist privacy amplification procedures not based on universal2 hashing, such as Trevisan’s extractor [27]. However, the qualitative properties of those procedures are mostly similar to those of the Leftover Hashing Lemma. Furthermore, as mentioned there, additional care is necessary when analyzing privacy amplification if the pre-amplification strings are not bitstrings of fixed-length, since this may not technically fit the conditions of the Leftover Hashing Lemma. We briefly discuss this in Section 5.6.2 later, and a full rigorous treatment of this issue can be found in Ref. [114].
@@ 4.6.1 Source maps
Formally, given the set of (infinite-dimensional) states $\{\rho_{i}\}$ used in the protocol, we find a set of (finite-dimensional) states $\{\tau_{i}\}$ such that there exists a channel $\Phi$ where $\rho_{i}=\Phi(\tau_{i})$ for all $i$. This channel $\Phi$ constitutes the source map which is ‘given’ to Eve, and thus any security analysis that assumes that Alice prepares the states $\{\tau_{i}\}$ would also hold for the actual protocol where Alice prepares the states $\{\rho_{i}\}$ as this only gives Eve more power. See [75, Lemma 8] for a more formal proof of this.
@@ 4.6.2 Squashing maps
Simple squashers [10, 40] Protocol-dependent construction — need to check for each protocol separately (e.g. - equal loss in all detectors for active BB84 setup, no dark counts.) Proof-technique independent [75, Lemma 6]
@@
Flag-state squasher [126] Block-diagonal structure of POVM elements EAT-based proof techniques [49, Appendix B], IID analysis [126, Theorem 1]
Dimension-reduction method [115] None EAT-based proof techniques [^29] Note that although this has been used for CV-QKD protocols, the extension to DV-QKD protocols is straightforward. [83, Theorem 2], IID analysis [115]
Weight-preserving flag-state squasher [75, Section IV.B.1.] Block-diagonal structure of POVM elements Proof-technique independent [75, Lemma 6]
Proof-technique independent squashers: The simple squashers and weight-preserving flag-state squasher use the same general principle [111, 10, 108, 40] depicted in Fig. 2, similar to source maps. These squashers allow us to directly replace the infinite-dimensional POVM with a finite-dimensional POVM for a QKD security proof. That is, they work by proving that the security of the QKD protocol with the finite-dimensional POVM implies the security of the QKD protocol with the infinite-dimensional POVM [75, Lemma 6]. The security proof for the former protocol can then be obtained via any proof technique of one’s choice.
@@ 5.1.2 Iterative sifting
Protocols with on-the-fly announcements (but with a fixed number of total rounds) are easily handled in the GEAT framework [67], but then suffer from the sequential attack assumption (see Section 6.4.2). For techniques to additionally handle the scenario where the total number of rounds also depends on these announcements, see Ref. [101] for methods in phase error correction (Section 6.2) and EUR (Section 6.1) frameworks.
@@ 5.2.1 Aborting during authentication
Moreover, in the setting where the authentication channel is allowed to abort, we note that typical constructions of such channels result in only the receiving party learning of the abort, while the sending party remains unaware. This asymmetry implies that Eve can always force an abort on one side while the other party accepts (by exactly mimicking the honest behaviour up until the last message sent between the parties, which she then interferes with to force the receiving party to abort). Note that this scenario is not covered by the security definition in Definition 2.1, i.e, Definition 2.1 cannot be satisfied for any QKD protocol where authentication is allowed to abort and only the receiving party learns of the abort.
However, we do not expect this gap to pose a fundamental problem. Specifically, we believe that QKD security proofs established under the assumption that authentication never aborts can be suitably adapted to the setting where the authentication may abort – potentially resulting in a final state where only one party has accepted (see also [82, Section VII A]). However, we note that to our knowledge, the correct security definition in this setting is only explicitly discussed in Ref. [89]. Moreover, the exact modifications needed to adapt the usual security proofs (established under the assumption that authentication never aborts) to this setting have not yet been discussed in the literature.
5.2.2 Aborting during error verification
@@ 5.3.1 Collective attacks are not a good approximation to practical restrictions
A common assumption made in some existing work is that Eve’s attack is independent and identically distributed (IID) in each round, i.e., she interacts with Alice’s signal states in each round exactly the same way, independent of prior protocol rounds. We emphasise here that this assumption cannot be motivated by limiting Eve to currently practical attacks. That is, with currently available technology, it is easy to implement a non-IID attack.
@@ 5.3.3 Use of concentration inequalities
Hoeffding’s Inequality [47, Theorem 2] $X_{1},\dots,X_{n}$ independent random variables with $a_{i}\leq X_{i}\leq b_{i}$. $\Pr[\absolutevalue{\sum_{i}(X_{i}-E[X_{i}])}\geq t]\leq 2\exp(-\frac{2t^{2}}{\sum_{i}^{n}\left(b_{i}-a_{i}\right)^{2}})$
@@
Serfling inequality [93, Corollary 1.1] $X_{1},\dots,X_{n}$ drawn randomly without replacement from values $x_{1},\dots,x_{N}$ for $n\leq N$ $\Pr[\sum_{i}X_{i}-n\mu\geq nt]\leq 2\exp(-\frac{2nt^{2}}{(1-f_{n}^{*})(b-a)})$, for $\mu=\sum_{i}x_{i}/n,f_{n}^{*}=\frac{n-1}{N},a=\min_{i}x_{i}$ and $b=\max_{i}x_{i}$.